Two angles are supplementary if the sum of their measures is 180 degrees. The measure of the first angle is 18 degrees less than two times the second angle. Find the measures of each angle. I need the whole steps please!!!!! i need this answer ASAP THANK U
Answer by Love math(28)

x + x = 180
(2x -18)+ x = 180
3x -18 =180
3x = 180+18
x = 198/3
x =66
second angle is 66
first angle is 180-66=114